OpenStudy (anonymous):

A Labrador leaps over a hurdle. The function f(t) represents the height of the Labrador above the ground, in inches, at t seconds: f(t) = -16t2 + 20t

OpenStudy (anonymous):

A foxhound jumps over the same hurdle. The table shows the height of the foxhound above the ground g(t), in inches, at t seconds: Time (t) g(t) 0 0 0.4 7.04 0.6 8.64 0.75 9 1.0 8 1.5 0 Part A: Compare and interpret the maximum of f(t) and g(t)? (4 points) Part B: Which function has a greater x-intercept? What do the x-intercepts of the graphs of f(t) and g(t) represent? (4 points) Part C: Determine the y-intercepts of both functions, and explain what this means in the context of the problem. (2 points)

OpenStudy (anonymous):

@amistre64 @iGreen @ganeshie8

OpenStudy (anonymous):

@Preetha @Michele_Laino @nincompoop

OpenStudy (anonymous):

i got part A The maximum for t is 1.5 and the maximum for g(t) is 9. What that means is the foxhound jumped 9inches above the ground and the maximum for t means how long it took him.

OpenStudy (michele_laino):

for the first function, namely f(t), we have to find the coordinates of the vertex of the parabola f(t)

OpenStudy (michele_laino):

the coordinate of the vertex are V: \[\Large V = \left( { - \frac{b}{{2a}}, - \frac{{{b^2} - 4ac}}{{4a}}} \right)\] where a = -16, b= 20,a nd c=0, since I'm referring to a generic parabola: f(t)= a*t^2+ b*t+c

OpenStudy (michele_laino):

and*

OpenStudy (anonymous):

okay so for part B: i think its the first equation

OpenStudy (anonymous):

i think its the first equation and the x represents how much time it took for him to leave and ground and get back to the ground

OpenStudy (michele_laino):

we have to write the quadratic function, using the coordinates from your table

OpenStudy (anonymous):

okay.

OpenStudy (michele_laino):

let's conjecture this function for g(t): \[g\left( t \right) = a{t^2} + bt\] I wrote that function, since, from your table, the point (0,0) has to check the unknown function g(t)

OpenStudy (michele_laino):

now, the point t=1, g=8 belongs to the graph of g(t), so we can write this equation: \[a + b = 8\] furthermore the point t=1.5=3/2, and g=0, belongs to the graph of g(t), so we can write: \[a\frac{9}{4} + b\frac{3}{2} = 0\]

OpenStudy (michele_laino):

so, we have to solve this algebraic system: \[\Large \left\{ \begin{gathered} a + b = 8 \hfill \\ a\frac{9}{4} + b\frac{3}{2} = 0 \hfill \\ \end{gathered} \right.\] what are a, and b?

OpenStudy (anonymous):

do i find the common factor of 4 and 3 and then add them?

OpenStudy (michele_laino):

you have to find the least common multiple between 4 and 2

OpenStudy (anonymous):

okay so its 4 and then 3 gos to 6 and 9+6 is 15

OpenStudy (michele_laino):

here are your steps: \[a\frac{9}{4} + b\frac{3}{2} = \frac{{9a + 3b \times 2}}{4} = ...?\]

OpenStudy (michele_laino):

what is \[{3b \times 2}=...?\]

OpenStudy (anonymous):

uhm idk do i multply 3x2?

OpenStudy (michele_laino):

yes!

OpenStudy (anonymous):

6

OpenStudy (michele_laino):

ok! so we have: \[a\frac{9}{4} + b\frac{3}{2} = \frac{{9a + 3b \times 2}}{4} = \frac{{9a + 6b}}{4}\]

OpenStudy (anonymous):

so thats my quadratic equation

OpenStudy (michele_laino):

no, it is your second equation of the system, namely: \[\left\{ \begin{gathered} a + b = 8 \hfill \\ \frac{{9a + 6b}}{4} = 0 \hfill \\ \end{gathered} \right.\]

OpenStudy (michele_laino):

or:\[\left\{ \begin{gathered} a + b = 8 \hfill \\ 9a + 6b = 0 \hfill \\ \end{gathered} \right.\]

OpenStudy (anonymous):

okay.

OpenStudy (michele_laino):

now I solve the first equation for b, and I get: b= 8 -a then I substitute into the first equation, so I get: \[\Large 9a + 6\left( {8 - a} \right) = 0\]

OpenStudy (michele_laino):

please solve it for a, what do you get?

OpenStudy (anonymous):

uhm do you subtract 6 from both sides and get 9a(8-a)=-6 then what

OpenStudy (michele_laino):

no, you have to compute the multiplication between 6 and 8-a

OpenStudy (anonymous):

okay so 6 times 8 and then its 6a so 9a(48-6a)=0

OpenStudy (michele_laino):

hit: \[9a + 6\left( {8 - a} \right) = 9a + 48 - 6a\]

OpenStudy (michele_laino):

hint*

OpenStudy (michele_laino):

ok! so we get this equation: \[9a + 48 - 6a = 0\] what is a?

OpenStudy (anonymous):

so now you have 3a+48=0 and subtract 48 to both sides and divide by 3 and get a=-16

OpenStudy (michele_laino):

ok! now we have to substitute that value, into the first equation, and then we have to find the value of b, like this way: \[ - 16 + b = 8\] what is b?

OpenStudy (anonymous):

at 16 to both sides and you get b=24

OpenStudy (michele_laino):

perfect, so the equation of g(t), is: \[\Large g\left( t \right) = - 16{t^2} + 24t\]

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

now what about answering part B?

OpenStudy (michele_laino):

the x-intercept of f(t), is given solving the subsequent equation: \[\Large - 16{t^2} + 20t = 0\] what is the non null root of that equation?

OpenStudy (anonymous):

what does that mean

OpenStudy (michele_laino):

I factor out 4t and I get: \[4t\left( { - 4t + 5} \right) = 0\]

OpenStudy (michele_laino):

now I apply the product canceling law, and I get these roots: t=0, and t=5/4, so the x-intercept of f(t) is t= 5/4

OpenStudy (anonymous):

alright

OpenStudy (michele_laino):

now we have to do the same using g(t), namely: \[ - 16{t^2} + 24t = 0\] what are the roots of that equation?

OpenStudy (anonymous):

you could factor our 8t?

OpenStudy (michele_laino):

that's right!

OpenStudy (michele_laino):

\[8t\left( { - 2t + 3} \right) = 0\]

OpenStudy (anonymous):

t=-4 and t=3/8? or 2.67?

OpenStudy (michele_laino):

using the product canceling law, I get: t=0 and t= 3/2

OpenStudy (anonymous):

ohhh

OpenStudy (michele_laino):

so the x-intercept of g(t) is 3/2, whereas the x-intercept of f(t) is 5/4

OpenStudy (michele_laino):

is 3/2 grater than 5/4, or is 5/4 greater than 3/2 ?

OpenStudy (anonymous):

i think 3/2 is greater

OpenStudy (michele_laino):

ok! since: \[\frac{3}{2} = \frac{6}{4} > \frac{5}{4}\]

OpenStudy (anonymous):

ohh 5/4

OpenStudy (michele_laino):

so the x-intercept of g(t) is greater that the x-intercept of f(t)

OpenStudy (anonymous):

so they represent time of how long he was off the ground right?

OpenStudy (michele_laino):

that's right!

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

Now what about part C?

OpenStudy (michele_laino):

the y-intercepts for both functions is y=0 ince we have the subsequent draw: